# Statistical Principles for Platform Trials

### v2

@@ 4.3 Effect of correlation induced by sharing control on $V^{\star}$
which depends on $\rho_{ij}=Corr(T^{\star}_{i},T^{\star}_{j})=\lambda_{i}\lambda_{j}$, where $\lambda_{i}=\sqrt{1+\frac{n_{SC}}{n_{i}}}$.
Each subset $N$ has its own $\rho_{ij}$ matrix, and thus $\lambda$ matrices, depending not only on the ratio between sample sizes in the arms of treatments with index in $N$ and in the shared control arm but also on other covariates such as the baseline measurement of each patient. In assessing stability of error rate control of compounds with apparent efficacy, we consider the observed treatment outcomes as well as the covariates as known, So we compute $\rho_{ij}=\lambda_{i}\lambda_{j},i\neq j,$ for the $N$ of compounds that show efficacy. For modeling-based analyses, such computations are possible.
A B C D
A 0.0475 0.0093 0.0083 0.0087
B 0.0093 0.0475 0.0092 0.0096
C 0.0083 0.0092 0.0475 0.0086
D 0.0087 0.0096 0.0086 0.0475
Table 3: Matrix of variances of $Z_{i}$ (diagonal) and covariances between $Z_{i}$ and $Z_{j}$ (off-diagonal) with ERpF set at $5\%\times 4=0.2$ for Regimens ABCD of the ALS Platform trial. The elements sum to 0.2974 for $Var(V^{\star})$, so $StdDev(V^{\star})$ = 0.5454, which is larger than the ERpF of 0.2.
Table 4 shows $StdDev(V^{\star})$ the variability of $V^{\star}$ increases with increasing $\rho_{ij}$.
@@ 4.4 Conditional assessment of FAR control
Principle 4.1 (The Conditionality Principle).
@@
In our Platform trials setting, let $T_{i},i=1,\ldots,k,$ be the statistics for comparing the $k$ new compounds against a shared control in a platform trial. For proof of concept, consider platform trials in therapeutic areas where $T_{i}$ can be pivoted to form a $100(1-\alpha)\%$ confidence interval for the primary efficacy endpoints44footnotetext: which excludes the score test version of the logRank test, so that
the pivoted $T_{i}$ has a null distribution free of the efficacy parameter. Neurodegenerative diseases (such as Alzheimer’s) and Type 2 Diabetes Mellitus (T2DM) are examples of such therapeutic areas. For simplicity of illustration, suppose $\hat{\mu}_{i}$ and $\hat{\mu}_{SC}$ are Normally distributed and the sample sizes are large enough so that $SE_{\hat{\mu}_{i}-\hat{\mu}_{SC}}$ in (4.2) can be considered essentially a constant.
Let $V^{\star}$ denote the number of new treatments in a platform trial with their confidence intervals for efficacy not covering their true parameter values, Then sample mean $\hat{\mu}_{SC}$ of the shared control under-performing or over-achieving affects the distribution of $V^{\star}$. Correlation $\hat{\rho}_{SC}$ among the test statistics associated with the new treatments (induced by sharing a control) affect the distribution of $V^{\star}$ as well.
4.4.1 Effect of sample mean of shared control on $V^{\star}$
(a) $\hat{\mu}_{SC}=0.257$
1 SE above $\mu_{SC}$
(b) $\hat{\mu}_{SC}=-0.257$
0.5 SE below $\mu_{SC}$
(c) $\hat{\mu}_{SC}=-0.514$
1 SE below $\mu_{SC}$
Figure 3: Probability distribution of $V$, the number of times a type I error is committed. We assume 4 treatment arms with an $\alpha$ level of 0.05, giving an ERpF of $0.05\times 5=0.25$, a total sample size of 800, a ratio of 4:1 for treatment arm to control sample size allocation gives an unconditional correlation of new treatment arm minus control estimates of 0.5. We additionally assume that the true shared control mean $\mu_{SC}$ and treatment arm means $\mu_{i}$ are equal to 0 (but this can easily be modified to other numbers). In Figure 3(a), the distribution of V is fairly similar to the baseline. Since the $\hat{\mu}_{SC}$ is higher, treatment arms have more difficulty gaining approval. On the other hand, in Figure 3(b) and 3(c), since our $\hat{\mu}_{SC}$ value is much lower than out $\mu_{i}$, erroneous approvals become much easier, and thus we have higher weight on higher values of $V$.
We consider 2 cases: (1) positive $\hat{\mu}_{SC}$ and (2) negative $\hat{\mu}_{SC}$. For sake of simplicity, we consider the cases where all treatment arms are non-efficacious, i.e. $\mu_{SC}=0$.
For cases with a positive $\hat{\mu}_{SC}$, the situation is of concern for sponsors, who want treatments to be approved. The more positive the $\hat{\mu}_{SC}$ treatment arms have a more difficult time getting approved, resulting in $V$ having more weight on lower numbers of erroneous approvals and thus translating to having a higher $V^{\star}$. For regulators, especially in cases where treatment arms are non-efficacious, however, this does not pose an issue as the chances of erroneously approving a treatment are low as can be as seen in Figure 3(a).
For the cases with negative $\hat{\mu}_{SC}$, the situation becomes of concern to regulators. It is much easier to approve treatment arms. Due to this, however, the issue of erroneous approvals becomes much more apparent. In fact, it could even result in almost all the treatment arms being erroneously approved if the $\hat{\mu}_{SC}$ was low enough. While pharmaceutical companies may desire this outcome, this is not desirable from a consumer or regulation standpoint as it leads to non-efficacious drugs getting erroneously approved (and thus increase $V^{\star}$).
Let us assume a simple case where we have 5 trials that are not efficacious. We adjust $\hat{\mu}_{SC}$ so that it is standardized by $SE_{SC}$. If we have a high $\hat{\mu}_{SC}$ 0 such that all treatment arms would not be approved, our $V^{\star}$ would equal 5. On the other hand, we consider a $\hat{\mu}_{SC}<0$ such that all treatment arms would be erroneously approved. This would result in a $V^{\star}$ of 5.
This problem can be built upon to address different scenarios to account for the sum of the treatment arm sample size $n_{SC}$ and control arm $i$ sample size $n_{i}$ and correlation $\rho=\lambda_{i}\lambda_{j}$, where $\lambda_{i}=\sqrt{1+\frac{n_{SC}}{n_{i}}}$. This allows for consideration of $SE_{\hat{\mu}_{i}-\hat{\mu}_{SC}}$ as well, which is dependent on the population standard deviation $\sigma$ and the sample sizes $n_{SC}$ and $n_{i}$.
In general, having a larger shared control sample mean $\hat{\mu}_{SC}$ results in new treatment arms having a more difficult time getting approved, thus resulting in a lower $V^{\star}$. This makes intuitive sense since our comparison of control for efficacy would be higher so we would need a higher treatment arm score to gain approval. In the same manner, having a lower $\hat{\mu}_{SC}$ results in easier approval, resulting in more confidence intervals for efficacy not covering the true parameter values (and thus a higher $V^{\star}$ value). This effect on $V^{\star}$ demonstrate the importance of its consideration when testing for efficacy.
@@ 5 A Confident Platform Random Sizer App
An app has been created to not only help demonstrate the effect of conditioning on $\hat{\mu}_{SC}$, but also provide information about the calculated unconditional correlation between $T_{i}^{\star}$ and $T_{j}^{\star}$ and . We use the HEALY ALS trial as an example case. While all four drugs tested in Regimens ABCD’s lack of efficacy could be due to either the compound lacking efficacy or an overachieving shared control, for the sake of this example, we will assume both. Each arm had approximately 120 patients in the treatment and approximately 40 in the control, maintaining a treatment-to-control sample size ratio of 3:1. We can set the app so the maximum number of treatments to be erroneously approved is 4 with a 120:40 (3:1) sample ratio. Based on this information, we can then input values into our sliders. Note that ALS has a population standard deviation of 6.5 (i.e. $\sigma=6.5)$ which is used to calculate the standard errors internally in the app (See App: User Slider Controls)
Figure 4: Although none of drugs in Regimens ABCD in HEALY ALS trial were determined as efficacious, if we had an underperforming shared control sample mean, i.e. $\hat{\mu}_{SC}<0$, there would be a chance drugs could be approved with ease as shown above with a standardized $\hat{\mu}_{SC}=-1.5$
App: User Slider Controls $\alpha$ for Each Rx Arm: Type I error rate for each treatment arm. Note that ERpF is $k\alpha$
Assume $\alpha=0.5$ for each treatment arm.
@@
We assume an overachieving $\hat{\mu}_{SC}$, i.e. $\hat{\mu}_{SC}>0$ Different shared control values would affect the number of erroneous approvals
Maximum Number of Rx Arms Erroneously Approved: $n$, the number of treatment arms approved when they are non-efficacious. Assumes $n=k$, i.e. all treatments are non-efficacious.
$k=4$ so set $n=4$
@@
[…]
[Uncaptioned image] Unconditional Correlation Between New Treatment and Control Estimates: correlation between $T_{i}^{\star}$ and $T_{j}^{\star}$
While not given, this information can be found via the sample size ratio below.
Sample Size Ratio: Treatment to control sample size ratio for sample allocation. Note that while the app allows for one to go higher, it is generally not advised to go higher than $k:1$.
textit3:1 treatment-to-control sample size ratio
App: Output Distribution of V: Gives the expected value, standard deviation, and the distribution of $V$, the number of treatment arms erroneously approved.
Since we assume an overachieving $\hat{\mu}_{SC}$, we see that the probability of erroneously approving a treatment arm is low
@@
The study utilizes as 120:40 ratio, so a ratio of 3. The allocation approximately matches that in the study
@@
By choosing the value for the sample size ratio, the unconditional correlation of the new treatment arm and control estimates gets calculated automatically for the user, as is the case for the ALS study. This is beneficial to regulators in particular, who would have more interest invested in controlling this to stabilize the distribution of the control sample mean. By having lower correlation (and thus lower randomization ratio), $SE_{\hat{\mu}_{SC}}$ is lowered as well, making the estimate for less variable.
While this example utilizes an existing study, this application could also demonstrate its usage when designing a study, specifically into considerations of stability of the estimates of shared control and (unconditional) correlation between the difference between $T_{i}^{\star}$ and $T_{j}^{\star}$. In particular, if regulators demand a certain correlation threshold to be met, sponsors are able to determine the highest sample ratio randomization that can achieve said threshold. The app can be used to find the sweet spot where all parties of interest can be satisfied.

### v3

@@ 3.1.1 Effect of correlation induced by sharing control on variability of $V^{*}$
which depends on $\rho_{ij}=Corr(T^{\star}_{i},T^{\star}_{j})=\lambda_{i}\lambda_{j}$, where $\lambda_{i}=(1+\frac{n_{SC}}{n_{i}})^{-1/2}$.
Each subset $N$ has its distinct $\rho_{ij}$ matrix, and thus $\lambda$ matrices, depending not only on the ratio of sample sizes between the treatment arms with index in $N$ and the shared control arm, but also on other covariates like baseline measurement of each patient. To assess the stability of error rate control for compounds apparently with no efficacy, we consider the observed treatment outcomes as well as the covariates as known. Consequently, we compute $\rho_{ij}=\lambda_{i}\lambda_{j},i\neq j,$ among the compounds in $N$ that show no efficacy. Such computations are feasible within model-based analytical frameworks.
Table 1 illustrates that the standard deviation of $V^{\star}$, denoted as $StdDev(V^{\star})$, increases with higher correlation coefficients $\rho_{ij}$. If the treatment arms were independent, $StdDev(V^{\star})$ exceeds the mean $E(V^{\star})$. Moreover, correlations among the test statistics, caused by sharing a control, further increase the variability of $V^{\star}$ (the number of incorrect approvals). For example, a common correlation of 0.5 among the test statistics raises $StdDev(V^{\star})$ from 0.69 to 1.16. This inherent variability of the actual error rate in a specific family of clinical trials is a significant issue that, we believe, warrants greater attention in practice, irrespective of whether the error rate is measured as ERpF or ERFw.
@@
To facilitate the assessment of the impact of correlations, analogous to how internet commerce tracks cookies, we propose maintaining an informational “cookie” for each treatment or regimen in a platform trial.
Control Sharing Matrix
Each treatment or regimen should have a “cookie” that records the exact time interval of its enrollment period. By identifying the overlap between the enrollment periods of different regimens, it becomes possible to compute the pairwise covariance and correlation for comparisons between the active treatment and the control.
$n$ A B C D
A 122 122 164 103
B 122 120 122 103
C 164 122 120 103
D 103 103 103 75
Table 2: Control Sharing Matrix (CSM) for the HEALEY ALS Platform Trial
A B C D
A 1 0.498 0.425 0.478
B 0.498 1 0.496 0.476
C 0.425 0.496 1 0.476
D 0.478 0.476 0.476 1
Table 3: Correlations matrix between $\hat{\mu}_{i}-\hat{\mu}_{SC_{ij}}$ and $\hat{\mu}_{j}-\hat{\mu}_{SC_{ij}}$
for the HEALEY ALS Platform Trial
A B C D
A 0.0475 0.0093 0.0083 0.0087
B 0.0093 0.0475 0.0092 0.0096
C 0.0083 0.0092 0.0475 0.0086
D 0.0087 0.0096 0.0086 0.0475
Table 4: Matrix of variances of $Z_{i}$ (diagonal) and covariances between $Z_{i}$ and $Z_{j}$ (off-diagonal) with ERpF set at $5\%\times 4=0.2$ for Regimens ABCD of the ALS Platform trial.
Table 2 presents the Control Sharing Matrix (CSM) for the HEALEY ALS Platform Trial, inferred from the results available on clinicaltrials.gov. For each regimen, the diagonal elements represent the number of patients randomized to the active treatment. The off-diagonal elements indicate the number of patients shared in the control group, to which each active treatment is compared, for each pair of regimens. We can then derive Table 3, which presents the correlation matrix (Details of the derivation are available in the supplementary materials). Note that as mentioned before, the pairwise correlation in Table 3 can also be expressed as $\hat{\rho}_{ij}=\lambda_{i}\lambda_{j}$ for all $i<j\in N$, where $\lambda_{i}=(1+\frac{n_{SC}}{n_{i}})^{-1/2}$. These $\lambda$ values can be readily calculated using existing software, making this notation preferable for its ease of extension to more complex structures. For the HEALEY ALS Platform Trial, the approximated $\lambda=$ (0.674, 0.722, 0.671, 0.690). Based on equation (3.2), we can obtain variance-covariance matrix for $Z$ (see Table 4). Note that the elements sum to 0.2974 for $Var(V^{\star})$, resulting in $StdDev(V^{\star})$ = 0.5454, which exceeds the ERpF of 0.2.
3.1.2 Effect of sample mean of shared control on variability of $V^{*}$
Assuming, without loss of generality, that $\mu_{i}=0$ for $i\in N$. According to equation (3.1), a positive (or negative) estimated mean $\hat{\mu}_{SC}$ of the shared control can make it harder (or easier) for new compounds to obtain approval. In this paper, a positive $\hat{\mu}_{SC}$ is referred to as an under-performing estimated mean of the shared control, while a negative $\hat{\mu}_{SC}$ is termed an over-achieving estimated mean of the shared control.
An under-performing $\hat{\mu}_{SC}$ presents significant concerns for sponsors seeking treatment approvals. The higher $\hat{\mu}_{SC}$ becomes, the more challenging it is for treatment arms to receive approval. This scenario leads to high probability of $V^{\star}$ being towards lower numbers of erroneous approvals, resulting in a lower variability of $V^{\star}$. However, for regulators, especially when the treatment arms are non-efficacious, this situation is less problematic as it reduces the likelihood of erroneously approving a treatment.
With over-achieving $\hat{\mu}_{SC}$, the situation becomes a concern for regulators, as it is much easier to approve treatment arms due to the lower $\hat{\mu}_{SC}$. This increased ease of approval heightens the risk of erroneous approvals. In extreme cases, if $\hat{\mu}_{SC}$ is sufficiently low, almost all treatment arms may be erroneously approved. While sponsors might favor such an outcome, it is undesirable from a consumer and regulatory perspective because it leads to the approval of non-efficacious drugs, thereby increasing the probability of higher $V^{\star}$ and therefore enhancing variability of $V^{\star}$.
These two cases are demonstrated in Figure (1). Assume $\mu_{i}=0$ for $i=1,2,3,4\in N$, and $\alpha=0.05$. Then, Figure (1) compares the true probability distribution of $V^{*}$ when $\hat{\mu}_{SC}={\mu}_{SC}=0$ (the dash line) with those where $\hat{\mu}_{SC}$ is either greater than or less than ${\mu}_{SC}=0$ (the solid line). Figure 1a indicates an essentially negligible change in the probability of obtaining $V^{*}=1$, and the probabilities of obtaining $V^{*}=2,3,4$ are nearly zero, suggesting minimal impact on the variability of $V^{*}$. However, as shown in Figures 1b and 1c, the likelihood of obtaining $V^{*}=1$ and $V^{*}=2$ increases as $\hat{\mu}_{SC}$ decreases, thereby increasing the variability of $V^{*}$.
(a) $\hat{\mu}_{SC}=0.46$
1 SE above $\mu_{SC}$
(b) $\hat{\mu}_{SC}=-0.23$
0.5 SE below $\mu_{SC}$
(c) $\hat{\mu}_{SC}=-0.46$
1 SE below $\mu_{SC}$
Figure 1: Probability distribution of $V^{\star}$. In this scenario, we have 5 treatment arms, each with an $\alpha$ level of 0.05, resulting in an ERpF of $0.05\times 5=0.25$. The total sample size is 1200, with a 5:1 ratio of treatment arm to control sample size allocation, leading to an unconditional correlation of 0.5 between estimates in differences between each new treatment arm and the control. For illustration purpose, we assume that the true shared control mean $\mu_{SC}$ and treatment arm means $\mu_{i}$ both equal to 0 (although these values can be adjusted). In Figure 1(a), the distribution of $V^{\star}$ closely mirrors the underlying true distribution. With a higher $\hat{\mu}_{SC}$, it becomes more difficult for treatment arms to gain approval, leading to a negligible probability of $V^{\star}>0$. Conversely, Figure 1(b) and 1(c) show scenarios where the $\hat{\mu}_{SC}$ is much lower than $\mu_{i}$, facilitating easier erroneous approvals. In particular, a substantial increase for $V^{\star}=1$ and $V^{\star}=2$ would not cause only an increase in $V^{\star}$, but also its variability.
3.1.3 Effect of mean and correlation induced by sharing control on variability of $V^{*}$
It is to be noted that while both the mean and the correlation induced by sharing control individually affect the distribution of $V^{\star}$, it is important to consider their combined effect on the distribution of $V^{\star}$. Specifically, the variance of $V^{\star}$ is noteworthy. We standardized $\hat{\mu}_{SC}$ by $SE_{SC}$ which we denote by $\hat{\mu}_{SC}^{\star}$, i.e.
$\hat{\mu}_{SC}^{\star}=\frac{\hat{\mu}_{SC}}{SE_{SC}}.$
As seen in Figure 2, for cases when $\hat{\mu}_{SC}^{\star}$ is greater than the treatment arm mean $\hat{\mu}_{i}$, the performance of the standard deviation of $V$ is lower and much more stable, which translates to a more stable $V^{\star}$ value. On the other hand, when $\hat{\mu}_{SC}^{\star}$ is less than $\hat{\mu}_{i}$, the behavior of the standard deviation of $V$ is higher and much more variable, which translates to a more variable $V^{\star}$. Due to this behavior of $V^{\star}$, it is beneficial to look at the whole behavior of the probability distribution of $V$ rather than single instances of $V$.
Figure 2: The behavior of standard deviation for $V^{\star}$ is heavily affected by correlation and $\hat{\mu}_{SC}$. Assume $\mu_{SC}=0$, i.e. it is non efficacious. As we decrease the value of our $\hat{\mu}_{SC}$ and thus $\hat{\mu}_{SC}^{\star}$, the performance of standard deviation of $V^{\star}$ becomes less stable.
$\hat{\mu}_{SC}^{\star}$ = -3
$\hat{\mu}_{SC}^{\star}$ = -1.5
$\hat{\mu}_{SC}^{\star}$ = 0
$\hat{\mu}_{SC}^{\star}$ = 1.5
$\hat{\mu}_{SC}^{\star}$ = 3
Figure 3: Snapshot of different standardized $\hat{\mu}_{SC}$ values on correlation’s effect on standard deviation where $\mu_{i}=0$. As can be seen, when the $\hat{\mu}_{SC}^{\star}$ is less than the treatment mean, the behavior is less simple whereas for $\hat{\mu}_{SC}^{\star}$ is greater than the treatment mean, the standard deviation of V decreases as correlation increases.
@@ 3.2 Conditional assessment of FAR control
Principle 3.1 (The Conditionality Principle).
@@ 4 A Confident Platform Random Sizer App
An app [^2] Can be found at https://eouya001.shinyapps.io/ConfidentPlatformRandomSizer/ has been developed to not only illustrate the impact of conditioning on $\hat{\mu}_{SC}$ for the distribution of $V^{*}$, but also to provide insights into the calculated standard errors and unconditional correlation between $T_{i}^{\star}$ and $T_{j}^{\star}$. We simulate a case study where we test the efficacy of five drugs. The total resources available allow for 600 people. With this setup, we can adjust the input values using the app’s sliders. Notably, ALS has a population standard deviation of 6.5 ($\sigma$=6.5), which the app uses internally to calculate standard errors of $\hat{\mu}_{SC},\hat{\mu}_{i}$ and $\hat{\mu}_{i}-\hat{\mu}_{SC}$, $i=1,\ldots,n\in N$, where $n$ is the maximum number of treatment arms erroneously approved. An example case in which sponsors should be concerned is presented in App: User Slider Controls.
App: User Slider Controls $\alpha$ for Each Rx Arm: ERFw $\alpha$ for each treatment arm. Note that ERpF is $n\alpha$
$\alpha=0.05$ for each treatment arm.
@@
An example of an overachieving $\hat{\mu}_{SC}$, i.e. $\hat{\mu}_{SC}>0$. Different shared control values would affect the number of erroneous approvals
Maximum Number of Rx Arms Erroneously Approved: $n$, the number of treatment arms, when they are non-efficacious.
n=5
@@
Total sample size = 600
Unconditional Correlation Between New Treatment and Control Estimates: Correlation between $T_{i}^{\star}$ and $T_{j}^{\star}$
$\rho_{ij}=0.5$
[Uncaptioned image] Sample Size Ratio: Treatment to control sample size ratio for sample allocation. Note that while the app allows for one to go higher, it is generally not advised to go higher than $k:1$.
5:1 treatment-to-control sample size ratio
By selecting the sample size ratio, the app automatically calculates the unconditional correlation between the new treatment arm and control estimates. This feature is particularly beneficial to regulators, who have a vested interest in managing this aspect to stabilize the distribution of $V^{*}$. A lower correlation achieved by reducing the sample size ratio also decreases the variation of $V^{*}$.
App: Output Distribution of $V^{*}$: Gives the expected value, standard deviation, and the probability distribution of $V^{*}$, the number of treatment arms erroneously approved.
With an overachieving $\hat{\mu}_{SC}$, we see that the probability of erroneously approving a treatment arm is low
@@
The study utilizes as 500:100 ratio, so a ratio of 5.
@@
The app can be instrumental in designing new studies. It facilitates considerations regarding the stability of the estimates from shared controls and the (unconditional) correlation between $T_{i}^{\star}$ and $T_{j}^{\star}$. Specifically, if regulators set a required correlation threshold, sponsors can use the app to determine the maximum sample ratio randomization that meets this criterion. The app thus serves as a valuable tool for identifying the optimal balance that satisfies all stakeholders involved.
Within a platform trial, the randomization ratio $n_{Rx}:n_{C}$ of the new compound to the shared control can vary. Ratios range from a low of $1:1$ resulting in a sample size ratio of the shared control to each new compound of $k:1$—to a high of $k:1$, where the sample size ratio reverses to $1:1$. Given a fixed total sample size for the platform trial, selecting the appropriate randomization ratio involves considering multiple factors.
Patients generally prefer a higher $n_{Rx}:n_{C}$ ratio, as it increases their chances of receiving a new treatment compound. For individual drug sponsors, the advantage of participating in a platform trial includes being able to pool control patients, which allows for a higher randomization ratio than the standard $1:1$ typically seen in standalone studies. This higher ratio can facilitate more accurate effect estimations for new compounds.
However, the perspective of the master protocol sponsor tends to favor a lower ratio, which increases the sample size of the shared control, reducing the likelihood of an overachieving control sample that complicates the demonstration of efficacy across new compounds.
From a regulatory standpoint, the stability of the ERpF control is influenced by both the estimates of the shared control’s sample means and the correlation among the pivotal statistics $T^{*}$. A lower randomization ratio enhances the sample size of the shared control, mitigating the risk of an under-performing control sample that could erroneously ease the approval of ineffective compounds. Additionally, a lower ratio reduces the correlation among pivotal statistics, thereby enhancing the stability of ERpF control.
For example, in the HEALEY ALS Platform Trial, none of the four drugs tested in Regimens ABCD demonstrated efficacy in the primary endpoint, ALSFRS-R. [^3] Active treatments for Regimen A were Zilucoplan, B was Verdiperstat, C was CNM-Au8, and D was Pridopidine. The lack of efficacy could be attributed to either the inherent ineffectiveness of the compounds, an overachieving shared control, or a combination of both. Conversely, as illustrated in App: Regulator Concerns, with an under-performed shared control sample mean had (i.e., $\hat{\mu}_{SC}<0$), there might have been a higher chance of the drugs being approved too readily, as exemplified by a $\hat{\mu}_{SC}^{\star}=-1.5$. Such a scenario would raise significant concerns for regulators.
In sum, choice of the randomization ratio $n_{Rx}:n_{C}$ should take all these considerations into account. Our app lets one find the sweet spot.
App: Regulators Concerns We now present a scenario using our app where sponsors, such as MASS General, need not be concerned, but which should capture the attention of regulators. The following inputs and outputs illustrate a hypothetical situation: $\alpha$ for Each Rx Arm: $0.05$ Standardized $\hat{\mu}_{SC}$: $-1.5$ Maximum Number of Rx Arms Erroneously Approved: $n=5$ Total Sample Size: 600 Unconditional Correlation Between New Treatments and Control estimates: 0.5 Sample Size Ratio: 5:1
[Uncaptioned image] The probability distribution of $V^{*}$ indicates that an under-performing standardized $\hat{\mu}_{SC}$ significantly increases the likelihood of erroneously approving at least one treatment, compared to its true efficacy. Regulators should be particularly concerned about this, as it suggests a non-negligible probability that ineffective drugs could mistakenly reach the market.
5 False Discovery Rate (FDR) in a regulatory setting
5.1 FDR would change policy
False Discovery Rate (FDR) control has been proposed as a method for controlling the Type I error rate across studies in Platform trials (Robertson et al., (2023)). FDR is defined as the expected proportion of incorrect rejections (rejecting a true null hypothesis) among all rejections, both correct and incorrect  [^4] Rejecting a false null hypothesis is considered “correct”, while rejecting a true null hypothesis is “incorrect”..
Let $H_{0i},i=1,\ldots,k$ represent the null hypotheses that each compound in a Platform trial with $k$ studies lacks efficacy, then
$=$ $\mathrm{E}\left[\frac{\mbox{number of in-efficacious compounds incorrectly approved }}{\mbox{total number of compounds approved, both efficacious and in-efficacious}}\right]$
While current regulatory practice of controlling FAR (which is Tukey’s per family error rate across studies) limits the number of in-efficacious compounds incorrectly approved to 5-per-100, controlling FDR in Platform trials would increase that number to beyond 5-per-100 as we will show below.
Recall $N$ represent the subset of index $\{1,\ldots,k\}$ for compounds that lack efficacy and $|N|=n$. For each $i\in N$, $Z_{i}$ is the indicator that the $i^{th}$ confidence interval erroneously shows efficacy with $P(Z_{i}=1)$, and
$V^{*}=\sum_{i\in N}Z_{i}.$
Then
$\dfrac{\mathrm{EFpF}}{n}$ $=$ $\dfrac{E(V^{*})}{n}$ (5.2)
$=$ $\dfrac{\sum_{i=1}^{n}P\left\{Z_{i}=1\right\}}{n}$ (5.3)
$=$ $\sum_{v=1}^{n}\dfrac{v}{n}P\{V^{*}=v\}$ (5.4)
$\mathrm{FDR}$ $=$ $\sum_{v=1}^{n}\dfrac{v}{v+n^{\mathrm{false}}}P\{V^{*}=v\}$ (5.5)
$\leq$ $\sum_{v=1}^{n}\dfrac{v}{0+n^{\mathrm{false}}}P\{V^{*}=v\}.$ (5.6)
How ERpF and FDR count the number of incorrect approvals $v$ differently is evident by comparing (5.5) with (5.4). FDR attenuates the number of incorrect rejection $v$ by the number of correct rejection of false nulls which we denote by $R$.
From (5.6), we see $n^{\mathrm{false}}$ being not small (relative to $v$) allows inflation of $P\left\{Z_{i}=1\right\}$ while keeping FDR small. For simplicity, suppose each Platform trial has $k=2$ studies only, and suppose the compound in Study $w$ is highly efficacious but the compound in Study $v$ is in-efficacious:
$FDR$ $=$ $E\left[\frac{V}{R}\right]$
$=$ $\frac{1}{2}\times 10\%=5\%$
Switching to FDR control would thus be a policy change: For every efficacious compound approved, regulators reward the industry with an additional fungible False Approval Rate (FAR) token [^5] We term the tokens fungible because any company can cash them in for any compound. toward approval of in-efficacious compounds.
Adopting FDR control would represent a significant policy shift. Under this system, for every efficacious compound approved, regulators would essentially reward the industry with an additional fungible False Approval Rate (FAR) token [^6] We describe the tokens as fungible because any company can use them to seek approval for any compound. that could be used toward the approval of inefficacious compounds. It is important to note that we find no explicit language in the Food, Drug, &\&& Cosmetic Act or the Kefauver-Harris Amendment suggesting that the number of efficacious compounds correctly approved affects the guarding of the number of in-efficacious compounds incorrectly approved.
5.2 Integrity of error rates against manipulation is desirable
Whereas the False Approval Rate (FAR) is only concerned with the rate at which in-efficacious compounds are incorrectly approved, clearly the issue with FDR is its denominator contains both correct and incorrect approvals of compounds. Similar to the concern in Section 6 of Finner and Roter (2001), we wonder whether that might subject FDR to manipulation as follows.
If FDR control were implemented across platform trials involving multiple drug companies. It seems plausible that a strategic company might hesitate to submit their best compound, preferring instead to let other companies take that risk. Similarly, in platform trials with a single sponsor, there could be a temptation to include highly efficacious compounds that have low commercial value. This strategy could potentially be used to facilitate the approval of less efficacious but more profitable drugs.